2. (one point) Show that for any three events A, B, and C with P(C) >0, P(A U B|C) = P(A|C) + P(BIC) – P(AN B|C)

How can I prove this?
Answer #1

[: P/A/8)= P(A1B) P(AUB) c) = PLAIC)+P (B/C ) - PLANBIC) L.H.S = P(AUBIC) = P(AUB) nC) P(B) P(c) pllanc) ulen)) P(c) 1. P(AUB